drawing balls from an urn. Anne randomly chooses a single ball from the urn shown here, x represents the color of the ball chosen. Construct a complete probability distribution for the random variable x.
3 red
5 yellow
2 blue

Answer by stanbon(75887)

drawing balls from an urn. Anne randomly chooses a single ball from the urn shown here, x represents the color of the ball chosen. Construct a complete probability distribution for the random variable x.
3 red
5 yellow
2 blue
------
Ans: P(red) = 3/10
P(yellow) = 5/10
P(blue) = 2/10
----------------------
Cheers,
Stan H.
-------------------